OpenStudy (anonymous):

2 Ag+NiO2+H20------->Ag2O+Ni(OH)2 name the element oxidized and name the element reduced

thomaster (thomaster):

Oke the reaction \(\sf 2Ag+NiO_2+H_2O\to Ag_2O+Ni(OH)_2\) When an element gets oxidized, it means it gaines electrons This also means that the oxidation number decrease because eletrons have a negative charge When an element gets reduced, it loses eletrons, ant the oxidation number increases. So we need to find out where the oxidation number decreases and increases. Oxygen is naturally 2- and hydrogen is naturally 1+ So the elements that get reduced/oxidized are Ag (silver) and Ni (nickel) We'll start with Ag, it is in elemental form, so the oxidation number is 0. after the arrow, Ag is in combination with oxygen. We know oxygen is 2- \(\sf Ag_2O\) The overall charge is 0 the O is 2- We have 2x Ag so the oxidation number of Ag has to be 1+ (because 2x1=2+) Now we know that the oxidation number in Ag increases, this means Ag gets reduced.

OpenStudy (anonymous):

im so lost!

thomaster (thomaster):

Where are you lost?

OpenStudy (anonymous):

everything :( just keep going im trying to keep up

OpenStudy (anonymous):

whats the element being oxidized

thomaster (thomaster):

I can help you understand this. But just telling which one gets oxidized won't help you very much. Lets do the same thing as above for nickel but this time in steps. Nickel is in combination with oxygen before the arrow. \(\Large\sf NiO_2\) We know that \(\sf O_2\) = 2- so what is the oxidation number for 2xO?

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

+4

thomaster (thomaster):

no -4 is correct :)

OpenStudy (anonymous):

okay now what?

thomaster (thomaster):

So when O2 is -4 what is the oxidation number of nickel?

OpenStudy (anonymous):

+4?

thomaster (thomaster):

Exactly! ;) Now we have to remember that Ni = 4+ We do the same thing for nickel on the other side of the arrow. \(\large\sf Ni(OH)_2\) O= 2- H= 1+ so what is the total oxidation number of \(\sf (OH)_2\) ?

OpenStudy (anonymous):

-2?

OpenStudy (anonymous):

2*-2 than add 2*1 and get -2...did i even do that right?

thomaster (thomaster):

\(\sf (OH)_2\) we have 2x O = 2x2- = 4- and 2x H = 2x1+ = 2+ -4+2 = -2 So you're right :) Now we know that (OH)2 = -2 what is the oxidation state of nickel?

OpenStudy (anonymous):

+2

OpenStudy (anonymous):

?

thomaster (thomaster):

Yes

OpenStudy (anonymous):

my online program is gonna lock me out soon so spending so much time on one problem can you just give me the answer please lol

thomaster (thomaster):

So before the arrow, nickel was 4+, and after the arrow its 2+ that means the oxidation number decreases, so nickel gets oxidized

thomaster (thomaster):

So the answers is: Ag = reduced Ni = oxidized

OpenStudy (anonymous):

i got it wrong, its the other way around :(

OpenStudy (anonymous):

i give up, lol thanks for your help though

thomaster (thomaster):

Was my answer wrong?

OpenStudy (anonymous):

it was wrong bc Ag=oxidized and Ni= reduced

thomaster (thomaster):

hmm I did something wrong here... lemme check it

thomaster (thomaster):

the oxidation number of nickel decreases, so it's reduced. I did something wrong at the silver.

OpenStudy (anonymous):

okay well thank you. i give up on my homework for now lol

thomaster (thomaster):

I see what went wrong I made a mistake at the oxidation number of Ag in AgO2 this is 2+, not 2- so the oxidation number of silver increase, so it's oxidized I'm terribly sorry for this :(

OpenStudy (anonymous):

your fine

thomaster (thomaster):

I must admit, this stuff can be pretty confusing sometimes :P